### v2

# Influence of finite-temperature effects on CMB power spectrum

Abstract
[…]
1 Introduction
Understanding the cosmological parameters that govern the Universe is central to modern cosmology [1, 2], with models of cosmological parameter estimation evolving to incorporate increasingly sophisticated methods and data sources. Among the most influential studies, the Planck 2018 results [3] have provided a comprehensive analysis of cosmological parameters derived from CMB anisotropies, utilizing high-precision data to set a benchmark for subsequent analyses. These measurements impose stringent constraints on key parameters such as the density of different components of the Universe, the Hubble constant, and the spectral index. As a result, they have guided our understanding of the Universe’s composition and evolution, and provided a solid foundation for both theoretical models and observational research. However, while these results have enhanced our ability to probe the cosmological landscape, the models themselves are still evolving. For instance, there are ongoing efforts to refine the assumptions underlying the cosmological constant (CC) and the potential roles that quantum effects may play over cosmic time. While the conventional hydrodynamic approach incorporates temperature via classical thermodynamics, it is crucial to also consider the temperature effects from finite-temperature (finite-T) quantum field theory (QFT) (see [4] for a review), as recently elaborated in references [5] [6], and [7]. Given the high temperatures in the early Universe, these effects are unlikely to be disregarded when studying its history. This work aims to analyze the implications of a finite-T-corrected cosmological parameters in light of these considerations.
The conventional hydrodynamic approach as well as the one based on Boltzmann’s equations is essentially a classical framework: it does not take into account the quantum effects introduced by finite-T QFT. [^1] One might argue that quantum effects are already incorporated through the use of field-theoretic free energy. However, this formulation does not provide an adequate framework for the problem at hand; see Appendix A for a detailed discussion. In addition to classical thermodynamics, finite-T QFT effects should play a significant role in the cosmological constant, feeding into it through perturbative quantum corrections. Ideally, a second-quantized description whose effective description is provided by the hydrodynamics should be employed. Imagine obtaining the quantum action, i.e., 1PI action. (Only the first several terms in the derivative expansion will matter.) The loop contributions to the CC depend on the temperature. This in turn implies that in the corresponding FRLW-type cosmology the CC comes to be time-dependent through the time-dependence of the temperature [7]. [^2] A time-dependent CC implies the following for the conservation laws [7]. (See [8] and the references therein for earlier related works.) With separate conservation for matter and radiation, one does not impose its separate conservation condition for the CC. Instead, the conservation equation of the total stress tensor follows from the Einstein equation. This bears an interesting implication for renormalization of the Newton’s constant. Finite-T effects were considered earlier [5][6] [^3] These works are finite-temperature extensions of earlier works on foliation-based gravity quantization [9], with one key element being the physical state condition [10] (see [11] for a similar idea). In [12], the gauge $K=0$ (more generally $K=K_{0}$), where $K$ denotes the trace of the second fundamental form, was employed as a crucial ingredient for quantization. A similar approach was later applied in Witten’s recent work [13]. to tackle the long-standing cosmological constant problem. See [14], [15], and [16] for an inspiring review of the CC problem and subsequent works. The CC problem was originally observed from the technical feature of onshell renormalization that the one-loop correction is enormously larger than the observed value of the CC, which necessitated a fine-tuned cancellation between the renormalized value of the CC and the one-loop correction. However, what has been noted in [5, 6, 7] (see also [17] and [18]) was that if one takes the renormalized mass to be on the order of the temperature - which is hinted by the fact that the presence of temperature yields a $T^{4}$-order CC term at the quantum level, the one-loop correction becomes small and this naturally allows one to take a small value of the renormalized value of the CC. It has been proposed that the CC problem may well be a matter of how to manage perturbation theory and finite renormalization, instead of a genuine problem. See also [14, 15] and [19] for ealier and more recent related discussions, respectively.
We consider a temperature-dependent CC (and its implications for the other cosmological parameters) for several reasons. First, the temperature effects are unavoidable due to perturbative quantum gravitational influences, making it impossible to omit unless negligible. Since we’re considering periods around or before recombination, these finite-T QFT effects are not negligible, though small, especially given the observed smallness of the present-day CC. Second, these effects exhibit the characteristics needed for early dark energy (see [23] for a review of early dark energy), making the temperature-dependent CC a form of early dark energy present throughout the Universe’s history. Their contribution is significant in the radiation-dominant era and around last scattering.
[…]
More specifically we consider the model that has the following eight parameters,
[…]
It is worth emphasizing that the present approach differs significantly from that of the 2018 Planck collaboration. The latter adheres to the standard Bayesian parameter inference framework: the observed $C_{\ell}$’s are treated as realizations of random variables, while the cosmological parameters are modeled as random variables endowed with prior distributions. This is a well-established Bayesian statistical paradigm. In contrast, the current ML-enhanced deterministic approach treats the cosmological parameters as input variables in a polynomial regression model. (See the introduction of section 3 for the reason and justification for adopting a different approach.) The coefficients of the polynomial - which relate these parameters to the angular power spectrum - are determined through regression. Subsequently, the values of the cosmological parameters that minimize the discrepancy with the 2018 Planck power spectrum curve are identified. Although we are employing frequentist framework, this is not a standard frequentist approach: the cosmological parameters are promoted to input variables in the regression setup, while the coefficients of the polynomial - rather than the cosmological parameters themselves - assume the role of parameters in the statistical sense. The resulting model may be viewed as a surrogate model that approximates the mapping from cosmological parameters to observables. We have more on this in the body.
Both 7- and 8 parameter models demonstrate exceptional accuracy, with the values of the coefficient of determination $R^{2}$ nearing 99.9 % for training- as well as testing- dataset. This metric, with mean squared error (MSE), highlights the model’s ability to accurately capture the underlying variance of the distance variable while maintaining low error rates. The results are then compared to the CLASS default model (6 parameters) and the Planck 2018 model (7 parameters).
The rest of the paper is organized as follows. In section 2, we begin by recalling the key results from the work [7] on the finite-temperature quantum gravity (QG) effects. In section 2.1, we adopt Weinberg’s approach [1], essentially reproducing the power spectrum plot derived therein, and compare it with the Planck 2018 results [3]. Although Weinberg’s formalism is not suitable for a detailed comparison with current data, it serves as a useful starting point. This approach aids in understanding the underlying physics at intermediate steps and also provides an analytic expression for the final CMB power spectrum, $\frac{\ell(\ell+1)}{2\pi}C_{\ell}$. Following this, we examine the temperature-corrected cosmological constant (CC). Using Mathematica, we compute the values of the cosmological parameters, modified by the presence of the time-dependent CC, that best fit the fiducial Planck 2018 results. Many of the steps from Chapters 6 and 7 of [1] can be carried over with minimal modifications. For clarity, we demonstrate the analysis by including $\Omega_{2}$ but omitting $\Omega_{3}$ to keep the Mathematica calculations manageable in terms of memory usage. In Section 3, we begin by highlighting how our approach differs significantly from that of Planck 2018, and we explain the rationale behind this distinction. For computation, we employ a quantum-modified version of the CLASS model [22], which includes the additional parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, and obtain the best fit for the 2018 Planck data [3]. Initially, we perform a brute-force parameter scan, revealing that the theoretical power spectrum is sensitive to both $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, along with the other six parameters. Given the inherent randomness in the subsequent sampling methods adopted, we validate these results using an independent method. Specifically, we apply various machine learning techniques, including random forest and others, to assess the predictive performance of both the 7-parameter and 8-parameter models in estimating the distance variable. We use quartic regression - which performed best among the methods we tested - to optimize the feature values and minimize the predicted distance. The inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model proves to be valuable, as it improves the accuracy of the predictions while maintaining strong generalizability. Finally, in section 4, we conclude with a summary and potential directions for future research. In Appendix A, we distinguish the present framework from quantum field-theoretic thermodynamics formulated through Helmholtz free energy. Specifically, the former is inherently offshell, while the latter is restricted to an onshell description. A summary of the statistical and machine learning terminology used in this work is provided in Appendix B.
2 Finite-T effects on cosmological parameters
Finite-temperature one-loop renormalization analysis of the Standard Model coupled with gravity has been recently carried out [7]. The coupling constants, in particular, the cosmological constant, are shifted by temperature-dependent - thus time-dependent - quantum corrections. In this section, as a warm-up, we study the implications of the presence of $\Omega_{\Lambda_{2}}$ for $(h,\;\omega_{b},\;\omega_{cdm},\;A_{s},\;n_{s},\;\tau_{reio})$ by adopting and extending Weinberg’s semi-analytic formalism. The reason for employing the Weinberg’s formalism prior to a more precise numerical analysis by CLASS is that the former has the analytic final expression for the CMB power spectrum, $\frac{\ell(\ell+1)}{2\pi}C_{\ell}$. The Weinberg’s approach captures physics of the processes involved. By considering both approaches one can grasp intuition as well as accuracy.
2.1 Angular power spectrum in Weinberg’s approach
We begin with a lightning review CMB angular power spectrum computation in [1]. There are two main components in the analysis therein: the fractional ionization $X$ and the power spectrum itself $C_{TT,l}^{S}$ given in (2) and (8) below, respectively. The subscripts $TT$ denote temperature-temperature correlation and $S$ scalar. The fractional ionization $X$ is defined as
$X=\frac{n_{e}}{n},\quad n=n_{p}+n_{H}.$ (1)
where $n_{e},n_{p},n_{H}$ denote the density of electrons, protons, hydrogens, respectively. It obeys
$\frac{dX}{dT}=\frac{n\alpha}{TH}\left(1+\frac{\tilde{\beta}}{\Gamma_{2s}+\frac{8\pi H}{\lambda_{\alpha}^{3}n(1-X)}}\right)^{-1}\left(X^{2}-\frac{1-X}{S}\right)$ (2)
where
$H$ $=$ $7.204\times 10^{-19}\,T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}\,T+h^{2}\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}+h^{2}\,\Omega_{M}},$
$\alpha$ $=$ $\frac{1.4377\times 10^{-10}\,T^{-0.6166}}{5.085\times 10^{-3}\,T^{0.53}+1}\;cm^{3}s^{-1},\quad\tilde{\beta}=2.4147\times 10^{15}cm^{-3}\;e^{-39474/T}\,T^{\frac{3}{2}}\,\alpha,$
$\Gamma_{2s}$ $=$ $8.22458\,s^{-1},\quad\lambda_{\alpha}=1215.682\times 10^{-8}cm,\quad S=1.747\times 10^{-22}\;e^{157894/T}\,T^{\frac{3}{2}}\Omega_{B}h^{2}.$
The $h^{2}\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}$ term in $H$ is small and can be disregarded [1]. We have used $\tilde{\beta}$, instead of $\beta$ in [1], to avoid confusion with the integration variable $\beta$ appearing in (8) below. This result is obtained with a commonly adopted assumption that the radiation is viewed to suddenly go from thermal equilibrium with matter to a free expansion at the red shift $z_{L}$. The quantatative definition of the corresponding temperature $T_{L}$ is based on the so-called visibility function associated with the fractional ionization. To calculate $T_{L}$, consider the opacity
${\cal O}=1-\exp\Big[-\int_{t(T)}^{t_{0}}c\sigma_{\cal T}n_{e}dt\Big]=1-\exp\Big[-c\sigma_{T}\int_{T_{0}}^{T}dT^{\prime}\frac{n_{e}(T^{\prime})}{H(T^{\prime})T^{\prime}}\Big]$ (4)
where $\sigma_{{\cal T}}=0.66525\times 10^{-24}\;cm^{2}$, and the visibility function (it is usually defined in terms of $z$, $g(z)$.)
$g(T)=\frac{d}{dT}\exp\Big[-c\sigma_{T}\int_{T_{0}}^{T}dT^{\prime}\frac{n_{e}(T^{\prime})}{H(T^{\prime})T^{\prime}}\Big]=\frac{d}{dT}\exp\Big[-c\sigma_{T}\int_{T_{0}}^{T}dT^{\prime}\frac{n(T^{\prime})X(T^{\prime})}{H(T^{\prime})T^{\prime}}\Big]$ (5)
where $n_{e}=nX$ was used in the second equality. The explicit form of $n$ is
$n=0.76\;\frac{3H_{0}^{2}\Omega_{B}}{8\pi Gm_{p}}\Big(\frac{T}{T_{\gamma 0}}\Big)^{3}$ (6)
The last scattering temperature $T_{L}$ is defined by
$\frac{dg}{dT}\Big|_{T=T_{L}}=0.$ (7)
As for $C_{TT,\ell}^{S}$, after extensive analysis it was shown in [1] that the power spectrum formula is given by:
$\frac{l(l+1)}{2\pi}C_{TT,l}^{S}$ $=$ $\frac{4\pi}{25}T_{0}^{2}N^{2}e^{-2\tau_{\small{{reion}}}}\int_{1}^{\infty}d\beta\left(\frac{\beta l}{l_{{\cal R}}}\right)^{n_{S}-1}$ (8)
$\hskip-65.04256pt\left\{\frac{3(\beta^{2}-1)^{\frac{1}{2}}}{\beta^{4}(1+R_{L})^{\frac{3}{2}}}e^{-2\beta^{2}l^{2}/l_{D}^{2}}{\cal S}^{2}(\beta l/l_{T})\sin^{2}\Big(\beta l/l_{H}+\Delta(\beta l/l_{T})\Big)\right$.
$\left.\hskip-108.405pt+\frac{1}{\beta^{2}(\beta^{2}-1)^{\frac{1}{2}}}\Big[3{\cal T}(\beta l/l_{T})R_{L}-(1+R_{L})^{-\frac{1}{4}}e^{-\beta^{2}l^{2}/l_{D}^{2}}{\cal S}(\beta l/l_{T})\cos\Big(\beta l/l_{H}+\Delta(\beta l/l_{T})\Big)\Big]^{2}\right\}.\;$
Above, $N$ is a normalization paramter and $\tau_{\small{{reion}}}$ denotes the optical depth of reionized plasma. By following [1] we set
$\frac{4\pi}{25}T_{0}^{2}N^{2}e^{-2\tau_{\small{{reion}}}}=519.7\,\mu K^{2}.$ (9)
$R_{L}$ is defined as the ratio of the background baryon and photon densities
$R_{L}\equiv\frac{3}{4}\frac{{\rho_{B}}}{{\rho}_{\gamma}}\Big|_{z=z_{L}}=\frac{3\Omega_{B}}{4\Omega_{\gamma}}\frac{1}{(1+z_{L})}.$ (10)
The transfer functions ${\cal S},{\cal T},\Delta$ were numerically determined and given by
${\cal S}(\kappa)$ $\equiv$ $\left(\frac{5^{1/2}(0.1657\kappa)^{6}+(0.5116\kappa)^{4}+(1.209\kappa)^{2}+1}{(0.1657\kappa)^{6}+(0.4249\kappa)^{4}+(0.9459\kappa)^{2}+1}\right)^{2}$
${\cal T}(\kappa)$ $\equiv$ $\frac{\log[(0.124\kappa)^{2}+1]}{(0.124\kappa)^{2}}\sqrt{\frac{(0.2197\kappa)^{6}+(0.4452\kappa)^{4}+(1.257\kappa)^{2}+1}{(0.3927\kappa)^{6}+(0.8568\kappa)^{4}+(1.606\kappa)^{2}+1}}$
$\Delta(\kappa)$ $\equiv$ $\left(\frac{(0.2578\kappa)^{6}+(0.5986\kappa)^{4}+(1.1547\kappa)^{2}}{(0.2204\kappa)^{8}+(0.4581\kappa)^{6}+(0.8707\kappa)^{4}+(1.723\kappa)^{2}+1}\right)^{\frac{1}{4}}.$ (11)
The definitions of $l_{T},l_{D},l_{H},l_{{\cal R}}$ are as follows:
$l_{T}\equiv\frac{d_{A}}{d_{T}},\quad l_{D}\equiv\frac{d_{A}}{d_{D}},\quad l_{H}\equiv\frac{d_{A}}{d_{H}},\quad l_{\cal R}\equiv(1+z_{L})k_{{\cal R}}{d_{A}}$ (12)
where the arbitrary pivot scale $k_{\cal R}$ is chose to be =0.05=0.05, and $d_{T}$, the acoustic horizon distance at last scattering $d_{H}$, and the angular diameter distance $d_{A}$ are given by
$d_{T}$ $\equiv$ $\kappa\frac{a_{L}}{q}=\frac{\sqrt{\Omega_{R}}}{(1+z_{L})H_{0}\Omega_{M}}$
$d_{H}$ $\equiv$ $\frac{2}{H_{0}\sqrt{3R_{L}\Omega_{M}}\;(1+z_{L})^{3/2}}\ln\Big(\frac{\sqrt{1+R_{L}}+\sqrt{R_{EQ}+R_{L}}}{1+\sqrt{R_{EQ}}}\Big)$
$d_{A}$ $\equiv$ $r_{L}a_{L}=\frac{1}{H_{0}}\frac{1}{1+z_{L}}\int_{(1+z_{L})^{-1}}^{1}\frac{dx}{\sqrt{\Omega_{\Lambda}x^{4}+\Omega_{M}x}}$ (13)
respectively. The symbols $\kappa,R_{EQ}$ are defined as
$\kappa\equiv\frac{\sqrt{2}\;q}{a_{EQ}H_{EQ}},\quad R_{EQ}=\frac{3}{4}\frac{{\rho_{B}}}{{\rho}_{\gamma}}\Big|_{z=z_{EQ}}.$ (14)
For damping distances one has,
$d_{D}^{2}=d_{\mbox{Landau}}^{2}+d_{\mbox{Silk}}^{2}$ (15)
where, for Landau damping,
$d_{\mbox{Landau}}^{2}$ $=$ $\frac{3\sigma^{2}t_{L}^{2}}{8T_{L}^{2}(1+R_{L})}$ (16)
where $\sigma$ denotes the standard deviation of ${\cal O}^{\prime}(T)$ when approximated as a Gaussian distribution, and
$d_{\mbox{Silk}}^{2}=\frac{cR_{L}^{2}}{6(1-Y)n_{B0}\,\sigma_{{\cal T}}\,H_{0}\,\sqrt{\Omega_{B}}\,R_{0}^{9/2}}\int_{0}^{R_{L}}\frac{R^{2}dR}{X\,(1+R)\sqrt{R_{EQ}+R}}\Big(\frac{16}{15}+\frac{R^{2}}{1+R}\Big)$ (17)
for Silk damping. The paramter $t_{L}$ in the Landau damping can be computed by the general formula for the time for an event with the red shift $z$
$t(z)=\frac{1}{H_{0}}\int_{0}^{\frac{1}{1+z}}\frac{dx}{x}\Big[\Omega_{\Lambda}+\Omega_{M}x^{-3}+\Omega_{R}x^{-4}\Big]^{-\frac{1}{2}}.$ (18)
Before delving into the finite-temperature corrections in the next section, we would like to briefly comment on some of the approximations made in the analysis of [1]. In addition to the approximation of treating recombination as nearly instantaneous, several other approximations and simplifications were made. For example, the integrated Sachs–Wolfe effect, which primarily affects small values of $\ell$, was neglected. Additionally, the tight coupling approximation was used for the early photon-baryon fluid. (The CLASS model, in contrast, accounts for these effects with greater accuracy.) Furthermore, although less significant, there are minor ”discrepancies” between the results obtained by Weinberg and those presented in the present work. For instance, consider the last two rows in Table 2.3 of [1]. When using the same data from that table, the values for the standard deviations turned out to be slightly different in our analysis. This indicates that there is a small program-dependence in the computation of sigma.
2.2 Finite-temperature modifications
Finite-temperature effects feed the CC through perturbative loops. The CC thus becomes temperature-dependent for a FLRW background. The resulting CC can be viewed as a form of early dark energy that has been present and significant up until around matter-radiation equality. One might mistakenly assume that finite-temperature quantum loop-induced contributions to the cosmological constant, such as those involving photons and neutrinos, are already included as part of the radiation contribution. However, this view overlooks the role of virtual particles in the cosmological constant. The contributions from Standard Model (SM) particles (and gravitons) arise from the vacuum energy of virtual particles, while radiation corresponds to the energy density of physical, onshell particles moving through the Universe. To be specific, let us consider a system described by Boltzmann kinetic theory, such as a photon gas, coupled to a scalar field - e.g., the Higgs field - on a curved spacetime background. The total action is given by
$S=S_{EH}+S_{sf}+S_{kin},$ (19)
where $S_{EH}$ denotes the Einstein-Hilbert action with an appropriately small CC (see the discussion below), $S_{sf}$ is the action for the scalar field, and $S_{kin}$ accounts for the kinetic system. The explicit form of $S_{kin}$ is known in specific cases (see, e.g., [24, 25]), such as for an ideal gas, but will not be required for the present discussion. The corresponding metric field equation contains the well-known expression for the stress-energy tensor:
$T^{\mu\nu}=\frac{1}{\sqrt{-g}}\sum_{n}\delta^{3}(x_{n}(t)-x)\,p_{n}^{\mu}(t)\,p_{n}^{\nu}(t)\,E_{n}(t).$ (20)
To convey the idea without unessential complications, we consider a fixed, non-dynamical background metric. Also, we focus on the $T^{4}$ scaling associated with $\Omega_{\Lambda_{2}}$. The standard treatment focuses solely on the classical thermodynamics of the kinetic system, which yields the familiar exact $T^{4}$ dependence for the energy density, as noted, for example, in footnote 1 of Section 3.1 in [1]. [^4] One might raise that quantum field-theoretic thermodynamics can be conducted by way of introducing free energy. We take up this issue in Appendix A.
In contrast, the present work amounts to examining quantum field-theoretic effects arising from the scalar field sector $S_{sf}$ (and other SM particles and graviton, which do not change the qualitative behavior) at finite temperature. Evidently, such a scalar field cannot be viewed as radiation due to their heavy mass, and they contribute to the CC as virtual particles, i.e., through vacuum bubble diagrams. Those contributions take a form of an infinite series in powers of the temperature, the leading term of which scales as $T^{4}$.
As mentioned earlier, the crux of our approach lies in the introduction of $\Omega_{\Lambda_{2}}$ (and $\Omega_{\Lambda_{3}}$). The analysis in section 3 shows that $\Omega_{\Lambda_{2}}$ is on the order of $10^{-8}$. We view these parameters as phenomenological. However, in section 3.3, we will discuss the (un)naturalness of the order of $\Omega_{\Lambda_{2}}$, specifically its value of $10^{-8}$, by introducing a different renormalization scheme that is better suited for this purpose.
In anticipation of the analysis in section 3.3, let us bring home the distinction between the zero-temperature contributions and the finite-temperature contributions. For the CC problem, the role of the temperature-dependent terms was more highlighted in [5], [6]. This difference is due to the slight change of the renormalization schemes. Let us recall the renormalization of the cosmological constant proposed in [5] [6], and [7]. It was demonstrated that the one-loop correction is two or three orders of magnitude smaller than the observed value. The essential idea can be explained by considering a gravity-scalar system To be specific, let us consider an Einstein-scalar system:
$S_{\text{ren}}$ $=$ $-\int d^{4}x\Bigg[\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\Bigg].$ (21)
The cosmological constant (CC) value depends, of course, on whether one adopts the complete-square form or the version without the constant term. (Whether to use the complete-square form or the traditional one is not central to the CC problem.) For the onshell potential value, one gets [29][4] (see Appendix A for more details)
$V=-\frac{1}{90}\pi^{2}T^{4}+\cdots$ (22)
where the ellipsis includes terms that depend on the mass of the scalar. Let us choose the renormalized mass value appropriately to deal with the CC problem: set the renormalized mass of the scalar particle so that it is on the order of the temperature. (See [26] and references therein for earlier works where a CC resolution proposal involves a novel renormalization scheme.) At the same time let it be sufficiently smaller than the temperature so that their contributions to the CC can be disregarded. With this renormalization scheme, the leading term in the potential is $\sim T^{4}$, as indicated above. Evaluated at the present temperature of the Universe, this implies that there is a order $10^{4-5}$ gap between the value above and the observed value of the CC, a (potentially) big improvement from the gap noted in the CC problem. As anticipated in [5], once one considers the Standard Model, its particle content should further reduce the gap. To see this, note that a fermionic field contributes
$\frac{7}{8}\frac{\pi^{2}}{90}T^{4}+\cdots$ (23)
to the potential. By going over the contributions of the SM particles, which provides an additional factor of around 120, the CC order discrepancy is now reduced from $10^{60}$ to $10^{2-3}$. [^5] The details of renormalization analysis can be quite subtle. For the zero-T component, the bosonic and fermionic contributions tend to cancel each other out due to their opposite signs. However, for the $T^{4}$ term, the situation is different: the one-loop bosonic and fermionic contributions carry a negative sign, so the fermionic contributions add to the bosonic ones. Given that there are more fermionic degrees of freedom than bosonic ones in the SM (coupled with gravity), the relative magnitudes of these contributions must be carefully balanced by adjusting the renormalized masses of the bosonic and fermionic fields to ensure the correct sign for the overall cosmological constant (CC). This balancing act can be helped by adjusting, if necessary, the finite parts of the renormalization parameters, often denoted as $c$’s in renormalization theory. These $c$’s play a crucial role in the proposed resolution of the CC problem. This observation was based on the one-loop correction, which has been shown to be two or three orders of magnitude smaller than the observed value, as just stated. However, it is possible to introduce a classical contribution to the cosmological constant (CC), and it is standard practice in renormalization to include such a classical term, the renormalized CC, $\Lambda_{ren}$, with the corresponding expression added to the density parameter. With this approach, the unnaturalness of adding or subtracting extremely large and small numbers is no longer present [5, 6, 7]. (This approach, however, comes with a trade-off, as reviewed in section 3.3.)
With the role of the leading T-dependent part of the vacuum energy in dealing with the CC problem understood, one can consider the finite-T contributions for the present purpose. (Here we are implicitly employing a new renormalization scheme more convenient for the task at hand; see section 3.3. As we will discuss there in more detail, the motivation for this is to avoid the unnaturalness of the order of $\Omega_{\Lambda_{2}}$, $10^{-8}$.) Let us now consider conducting cosmological perturbation analysis with the 1PI action (with only the first few leading-order terms kept) coupled with the hydrodynamic (or Boltzmann) system. One of the implications of [5, 6, 7] is that the leading quantum corrections yield the following form of the time-dependent vacuum energy [^6] Since higher-order loop corrections are among the terms represented by $(\cdots)$, additional time-dependent terms with different $a$-scalings will arise. These terms are not considered in the present work, although they will be addressed in the conclusion.
$\Lambda(t)=\Lambda_{1}+\Lambda_{2}\left(\frac{a_{0}}{a}\right)^{4}+\Lambda_{3}\left(\frac{a_{0}}{a}\right)^{2}+\cdots$ (24)
where $\Lambda_{1},\Lambda_{2}$ are time-independent constants, with $\Lambda_{1}+\Lambda_{2}+\Lambda_{3}+\cdots$ representing the known present value of the CC. The time-dependent CC in eq. (24) introduces additional density parameters, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, corresponding to the $a^{-4}$- and $a^{-2}$-scalings, respectively, in addition to the usual density parameter $\Omega_{\Lambda_{1}}$. Although the CC term associated with $\Omega_{\Lambda_{2}}$ scales as $1/a^{4}$, it does not correspond to radiation, as already discussed. From another angle, note that radiation must satisfy $p=\rho/3$. Similarly, $\Omega_{\Lambda_{3}}$, which appears in the off-shell action through quantum corrections, is conceptually distinct from $\Omega_{K}$, which is associated with the parameter $K$ in the on-shell solution. [^7] At the technical level of the present approximate perturbation analysis, however, these two cannot be distinguished. In the analysis presented in section 3, we let $\Omega_{\Lambda_{3}}$ to take on the role of $\Omega_{K}$, setting the latter to zero in the modified CLASS system.
For Mathematica analysis we only include $\Omega_{2}$, but not $\Omega_{3}$, to keep the analysis manageable. The paramter $\Omega_{3}$ will be included in section 3 where the analysis by employing CLASS is carried out. The presence of an additional density parameter, $\Omega_{\Lambda_{2}}$, manifests itself only through the Hubble parameter $H$. Let us enumerate the channels through which the additional density parameter, $\Omega_{\Lambda_{2}}$, modifies various cosmological parameters. First, the presence of $\Omega_{\Lambda_{2}}$ modifies the fractional ionization $X$ through the modified $H$,
$H=7.204\times 10^{-19}\,T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}\,T+\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}+\frac{T}{2.725}\Omega_{\Lambda_{2}}h^{2}}.$ (25)
The modification of X leads, via eq. (7), to a modified value of $z_{L}$, thus other quantities, such as $R_{L}$, as well. The power spectrum formula eq. (8) was obtained by first obtaining a semi-analytic solution - the analysis carried out in Ch. 6 of [1] - and substituting it into a two-point correlator of the temperature variation. The analysis of Ch. 6 of [1] can be carried over with the minor change of $H$, eq. (25), whenever it appears. As for $C_{TT,l}^{S}$ in eq. (8), one can use ${\cal T,S},\;\Delta$ without modifications. The finite-T corrections modifies $d_{A}$ and $d_{D},d_{T},d_{H},l_{\cal R}$ as well through the modification of the former:
$d_{A}=\frac{2.9979\times 10^{3}}{(1+z_{L})\sqrt{\Omega_{\Lambda_{1}}h^{2}x^{4}+\Omega_{M}h^{2}x+(\Omega_{\Lambda_{2}}+\Omega_{R})h^{2}}}.$ (26)
Fig. 1 shows sample plots of $X$ for different values of $\Omega_{\Lambda_{2}}$.
Figure 1: typical behaviors of the fractional ionization $X$ as a function of $\Omega_{\Lambda_{2}}$; four arbitrary values of $\Omega_{\Lambda_{2}}$ are chosen for demonstration
Although $X$ decreases slower toward lower temperature with increased values of $\Omega_{\Lambda_{2}}$, explicit analysis based on eq. (7) reveals that the finite-T effect actually increase the value of $T_{L}$.
(a)                    (b)
Figure 2: (a) $(h,\Omega_{B},\Omega_{\Lambda_{2}})$-interpolation    (b) $(h,\Omega_{M},\Omega_{\Lambda_{2}})$-interpolation; for (a) $\Omega_{M}=0.13299/h^{2}$. The best fit parameters are $(h,\Omega_{B},\Omega_{\Lambda_{2}},n_{s})=(0.765515,0.057172,0.00005,0.985064)$; for (b), $\Omega_{B}=0.02238/h^{2}$. The best fit parameters are $(h,\Omega_{M},\Omega_{\Lambda_{2}},n_{s})=(0.649838,0.272761,-1.59513\times 10^{-6},0.99)$
To obtain the CMB power spectrum in terms of the seven parameters—six usual ones,
$h,\Omega_{B},\Omega_{M},N,n_{S},\tau_{\small{{reio}}},$ (27)
plus $\Omega_{\Lambda_{2}}$—one must first derive the various parameters that appear in the angular power spectrum in eq. 8, such as $l_{T},l_{D},l_{H},l_{{\cal R}}$, as functions of the seven parameters. To keep the Mathematica numerical study manageable, we adopt the values of $N$ and $\tau_{\text{reio}}$ from [1] and focus on how the parameters depend on four cosmological parameters, $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$. The determination of the optimal $n_{s}$ can be done at the final stage. In other words, $n_{s}$ can be addressed after evaluating $C_{TT,l}^{S}$, once the dependencies of $l_{T},l_{D},l_{H},l_{{\cal R}}$ on the independent cosmological parameters are determined. Our main effort focuses on determining the dependencies of $l_{T},l_{D},l_{H},l_{{\cal R}}$ on the four parameters $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$.
The expressions for $l_{T},l_{D},l_{H},l_{{\cal R}}$ themselves involve parameters such as $z_{L},\sigma,R_{L},R_{EQ},t_{L}$ (where $\sigma$ denotes the standard deviation of ${\cal O}^{\prime}(T)$, where $\cal O$} is given in eq. (4), when approximated as a Gaussian distribution). These parameters need to be expressed in terms of $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$ in order to quantitatively determine the values of the four parameters for the best fit.
Although it would be ideal to express these parameters in terms of the full set $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$, this turns out to be a highly memory-intensive process that requires more powerful computational resources (as was the case for $\sigma$, for instance). We will not pursue this full task in the present work. Instead, we group the parameters into two sets: $(h,\Omega_{B},\Omega_{\Lambda_{2}})$ and $(h,\Omega_{M},\Omega_{\Lambda_{2}})$, and separately study the dependence of the various parameters on each of these two groups. For each set, $(h,\Omega_{M},\Omega_{\Lambda_{2}})$ and $(h,\Omega_{B},\Omega_{\Lambda_{2}})$, we compute the power spectrum. Two examples of the resulting plots are shown in Fig. 2.888A minor technical note: the solid line in the plot is based on the CLASS default.ini file, not the Planck 2018 data. As we will see in the next section, the two are visually indistinguishable. We used the CLASS data because its $\ell$-range is larger, while the Planck 2018 data extends only up to $\ell=2508$.
Weinberg’s approach is insightful and valuable for understanding the physics of key events during the intermediate period between the generation of the CMB and its observation today. However, as noted in his book, its primary aim is not the precise calculation of the power spectrum curve. For such purposes, computational tools like CAMB or CLASS are typically employed.
3 Modified CLASS and machine learning techniques
In this section, we detail our approach to parameter inference, which fundamentally diverges from the standard Bayesian methods employed by collaborations like Planck (2018). Our methodology emphasizes model building over rigorous parameter inference and focuses on constructing a phenomenological model with additional parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ that can accurately reproduce the observed cosmic microwave background (CMB) angular power spectrum. We perform two complementary analyses: a brute-force parameter scan and a subsequent analysis leveraging statistical and machine-learning techniques. A core component of our approach is the use of a finite-temperature modified version of the cosmological code CLASS. This allows us to perform almost the entire parameter exploration within its Python wrapper, CLASSy, significantly enhancing computational efficiency.
The primary motivation for our non-standard approach is the behavior of the newly introduced parameters, specifically $\Omega_{\Lambda_{2}}$. Initial theoretical estimates for its magnitude were found to be several orders of magnitude larger than the values that ultimately provided the best fit to the Planck 2018 data. Using these initial estimates resulted in power spectra that deviated significantly from the observed data. Instead of a standard Bayesian inference, which would be computationally prohibitive due to the vast parameter space and the lack of a reliable prior for $\Omega_{\Lambda_{2}}$, we adopted an iterative, trial-and-error procedure. This process involved gradually reducing the value of $\Omega_{\Lambda_{2}}$ and performing extensive parameter scans until a good fit was achieved. Although this method lacks statistical rigor (see the conclusion for further discussion), it proved to be an effective strategy for finding the correct order of magnitude for this parameter.
For our analysis, we used the unbinned theoretical comparison spectrum file from the Planck 2018 collaboration:
COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt.
[…]
A key advantage of our methodology is its computational efficiency. The entire parameter search process is performed within CLASSy, with only a small, nearly instantaneous external code required to calculate and sort distance values. In contrast, a standard Bayesian analysis would necessitate additional software packages like MontePython or Cobaya. Our tests showed that integrating these packages can increase runtime by a factor of six to eight. Given that our data collection alone took over three months, this efficiency was critical. While a standard Bayesian approach might require a supercomputer for a similar task, our method provides a more accessible and rapid way to explore this complex parameter space. Once our trial-and-error method establishes a rough order of magnitude for $\Omega_{\Lambda_{2}}$, a more focused scan using tools like Cobaya becomes a feasible and less time-consuming next step.
In the modified version of CLASS [22], the original Hubble constant $H(t)$ is replaced by the following that contains the finite-T corrections:
$H=7.204\times 10^{-19}\,T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}\,T+\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}+\frac{T}{2.725}\Omega_{\Lambda_{2}}h^{2}+\frac{2.725}{T}\Omega_{\Lambda_{3}}h^{2}}.$
Once the values of the cosmological parameters are fixed, the power spectra are computed using the modified CLASS. A large number of parameter sets can be scanned by running the CLASS Python wrapper, CLASSy. Each resulting spectrum is then fed into a code that calculates the Euclidean distance between the generated curve and the Planck data [3], COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt. By these comprehensive scans and distance calculations we obtain tens of millions of data points. The details of the brute-force scan will vary slightly each time because the data collection procedure involves random sampling. It will be nice to check the qualitative features of the results by an independent method. For this purpose we employ machine learning techniques and evaluate the predictive performance of the models in estimating the distance.
[…]
The objective is to determine which model best explains the variance in the target variable, the distance $d$, while balancing model complexity and performance. Model selection is based on maximizing $R^{2}$ and minimizing the MSE in both training and testing datasets. Using quartic regression - which proved to be the best-performing among the models we tested, [^10] Regression likely performs well because the distance can, in principle, be expressed as an analytic function of the cosmological parameters. The regression model effectively approximates its multivariable Taylor expansion. including random forest and others - we compare the accuracy and generalization capability of both approaches. The comparison shows that including the additional parameters from finite-temperature QG effects modifies the values of the cosmological parameters.
Variable Count 7-Variable Model ($\Omega_{K}$, 6 variables) 8-Variable Model ($\Omega_{\Lambda_{2}}$, $\Omega_{\Lambda_{3}}$, 6 variables)
Mean STD Range Mean STD Range
distance 5000 55.75 8.19 [28.12, 66.92] 58.80 9.70 [26.83, 71.76]
$\Omega_{\Lambda_{2}}$ 5000 -2.07E-08 2.32E-11 [-2.075E-08, 2.068E-08]
$\Omega_{\Lambda_{3}}$ 5000 0.00233 0.000003 [0.00233, 0.002342] 0.004289 0.00002 [0.00425, 0.00431]
$h$ 5000 0.676876 0.000624 [0.676, 0.678] 0.675491 0.000325 [0.675, 0.676]
$\omega_{b}$ 5000 0.022376 0.000027 [0.02232, 0.02242] 0.022393 0.000028 [0.02235, 0.02245]
$\omega_{cdm}$ 5000 0.119924 0.000207 [0.11944, 0.12033] 0.119853 0.00018 [0.11944, 0.12033]
$A_{s}$ 5000 2.10E-09 3.58E-12 [2.09E-09, 2.1056E-09] 2.10E-09 2.98E-12 [2.0967E-09, 2.1100E-09]
$n_{s}$ 5000 0.965761 0.000491 [0.965, 0.967] 0.965896 0.000487 [0.965, 0.967]
$\tau_{reio}$ 5000 0.053948 0.000757 [0.052, 0.055] 0.054819 0.000591 [0.054, 0.056]
Table 1: Statistical summary of variables in the 7-parameter and 8-parameter models
3.1 Brute-force scan by CLASSy
In the analysis by employing CLASS, we first conduct brute-force scans to systematically explore the parameter space for cosmological models. By generating discrete power spectra over the multipole moment parameter $\ell$ for suitable ranges of the parameters, their distances from the Planck 2018 experimental curve can be calculated. This approach requires extensive scanning of the parameter space, ultimately determining the set of cosmological parameters that minimizes the distance to the Planck 2018 curve.
As stated in the beginning, the adoption of a distance-based parameter fit is partly motivated by the challenge of determining an appropriate range for $\Omega_{\Lambda_{2}}$. Given specific numerical values for $\Omega_{\Lambda_{2}}$ (along with other parameters), one can efficiently and deterministically compute the distance to the Planck power spectrum curve. This contrasts with the Planck analysis framework - such as that implemented via MontePython - which typically relies on stochastic sampling methods (e.g., MCMC), requiring significantly more computational effort for parameter space exploration. Furthermore, to enhance the robustness and efficiency of our analysis, we incorporate machine learning techniques to guide and refine the parameter search, enabling a more targeted and accelerated convergence toward the best-fit regions.
Using CLASS, specifically its Python wrapper CLASSy, we conduct extensive parameter scans, ultimately collecting several tens of millions of datasets. This immense volume of data is necessary to perform various checks and validate the robustness of the results. Based on these preliminary explorations, appropriate parameter ranges are carefully refined. From these final intervals, we collect 400,000 datasets for the 7-parameter case and 300,000 datasets for the 8-parameter case. The number of the datasets for the 8-parameter model was deliberately chosen less than that of the 7-parameter model. This was to bring out the fact that the former performs better in spite of this disadvantage. See the related comments below.
The impact of inclusion of $\Omega_{\Lambda_{2}}$ is reflected on the actual minimum values of the distance and the corresponding parameters set. For the 7-parameter model, the minimum distance of the 400,000 datasets is [^11] To provide a sense of the accuracy associated with a distance value around 30, consider the following example. Using the CLASS configuration file base_2018_plikHM_TTTEEE_lowl_lowE_lensing.ini, and changing the maximum multipole $\ell$ from 2500 to 2508, yields a computed distance of 56.1 (rounded) when compared to the reference data in COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt. Since base_2018_plikHM_TTTEEE_lowl_lowE_lensing.ini is designed to reproduce the Planck 2018 baseline results, the CLASS-generated spectrum should closely match the Planck theoretical spectrum. The observed discrepancy may arise from factors such as numerical precision, CLASS version, or calibration differences.
$d=28.122388$ (29)
which occurs at
$\Omega_{K}$ $h$ $\omega_{b}$ $\omega_{cdm}$ $A_{s}$ $n_{s}$ $\tau_{reio}$
0.002339 0.677333 0.022387 0.119889 $2.098889\times 10^{-9}$ 0.9660 0.054000
whereas for the 8-parameter model, the minimum of the 300,000 datasets is
$d=26.832229$ (30)
with
$\Omega_{\Lambda_{2}}$ $\Omega_{\Lambda_{3}}$ $h$ $\omega_{b}$ $\omega_{\text{cdm}}$ $A_{s}$ $n_{s}$ $\tau_{\text{reio}}$ $2.073444\times 10^{-8}$ 0.004297 0.675500 0.022394 0.119889 $2.101111\times 10^{-9}$ 0.966000 0.054444
These minima are actual data values. Note the negative sign of $\Omega_{\Lambda_{2}}$ to which we will come back later. Below we obtain the regression functions of these two datasets and obtain the minima predicted by the functions. Those minima will in turn be checked against the values (given in (33)) yielded by directly running CLASS with the corresponding values of the parameters as input. To provide some context for the distance values, we show two plots in Fig. 3. As seen in these plots, the curves with distances in the range of several hundreds are nearly identical to the Planck 2018 curve, appearing directly on top of it.
(a)                    (b)
Figure 3: The dotted plot in panel (a) represents a curve with a distance $d\approx 1200$ from the Planck 2018 data, while the dotted plot in panel (b) corresponds to a distance $d\approx 3400$.
Given that the 8-parameter model includes one additional parameter, a fair comparison would require a dataset larger than 400,000 by some appropriate factor to adequately cover the expanded parameter space. Despite this discrepancy, the results indicate that the minimum distance to the Planck 2018 curve is smaller for the 8-parameter model, an important feature shared by the statistical analysis below.
3.2 Analysis by statistical and machine learning techniques
Involving random sampling, the specifics of the brute-force scan may vary slightly with each iteration. It would be reassuring to access the qualitative aspects of the results using an independent method. For this, we apply statistical and machine learning techniques to evaluate the predictive performance of the two cosmological models. The minimum distance predicted by each model is explicitly checked by running CLASS with the corresponding set of the cosmological parameters; see eq. (33) below. Here are some details of the analysis.
The dataset comprises 5,000 observations, i.e., sets of the parameters, for both the 7-parameter and 8-parameter models. To ensure sufficient data for model validation using 10-fold cross-validation, with each fold containing at least 500 samples, the 5,000 minimum-distance data samples are selected from a total of 400,000 and 300,000 CLASSy-generated datasets for the 7-parameter and 8-parameter models, respectively. Each dataset is then split into a training set (70%), consisting of 3,500 datasets, and a testing set (30%), consisting of 1,500 datasets. To address differences in the scales of independent parameters, all features are standardized, which transforms each variable to have a mean of 0 and a standard deviation of 1. This mitigates bias toward features with larger ranges and ensures uniform contribution of all variables to the regression analysis.
The performance of the polynomial regression models is evaluated using $R^{2}$, which quantifies the proportion of variance in the dependent variable modeled by the independent variables. Additionally, the MSE used to measure the average squared difference between predicted and actual values, providing a complementary metric to evaluate model performance. Both models achieve exceptional accuracy, with the $R^{2}$ values nearing 99.9%. The inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces additional complexity, which increases the data requirements for achieving stabilization on the learning curve. However, this added complexity ultimately enhances the model’s predictive accuracy. Table 1 presents a detailed statistical summary of the key variables in the 7-parameter and 8-parameter models. Notably, the addition of $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces subtle shifts in other variables, such as $h,n_{s},$ and $\tau_{reio}$. These shifts are accompanied by slightly increased variability in the distance value, as reflected in its higher standard deviation and wider range in the 8-parameter model compared to the 7-parameter model. A series of regression models, ranging from linear regression to polynomial regression up to degree 10, are evaluated for both datasets. The models are built and assessed separately for the 7-parameter and 8-parameter datasets. The quartic regression model (polynomial degree = 4) emerged as the best-performing one based on $R^{2}$ and MSE, as it consistently explains the largest proportion of variance in both the training and testing datasets, compared with higher-degree polynomials that exhibit overfitting tendencies.
To further assess the robustness and reliability of our regression-based parameter estimates, we apply a non-parametric statistical technique - bootstrap resampling - as a complementary approach to standard inference methods. While our methodology is primarily a model-building exercise rather than a rigorous parameter inference framework, bootstrap analysis offers an empirical means to quantify uncertainty in the optimized parameters and predicted outcomes. From the resulting distributions, we compute empirical 95% confidence intervals (CI) using the 2.5th and 97.5th percentiles, providing a data-driven estimate of sampling variability. These CIs, though not Bayesian credible intervals, serve as practical indicators of the stability and precision of our regression outputs.
Figure 4: 7-parameter model estimation and residuals
Figure 5: 8-parameter model estimation and residuals
3.2.1 Single run results
The $R^{2}$ metric is calculated separately for the training and testing datasets to evaluate the models’ ability to generalize. For the 7-parameter model, the quartic regression achieves an $R^{2}$ of 99.90% on the training set and 99.88% on the testing set, with MSE values of 0.0713 and 0.0795, respectively. In the 8-variable model, including $\Omega_{\Lambda_{2}}$ improved the training $R^{2}$ to 99.92%, though the testing $R^{2}$ remained at 99.88%, with MSE values of 0.0792 and 0.1096. While the addition of $\Omega_{\Lambda_{2}}$ slightly increases the error on the testing set, overall performance remained strong. Figs. 4 and 5 visually confirm these results, showing the predicted versus actual values and residual distributions for both models. The plots reveal a near-perfect alignment along the diagonal.
Figure 6: Learning curve of 7 parameter model
Figure 7: Learning curve of 8-parameter model
The learning curves for the 7-parameter and 8-parameter models, shown in Figs. 6 and 7, provide valuable insights into how training and validation accuracy evolve as the number of training samples increases. Both models ultimately achieve high accuracy, with training and validation accuracies converging near 99%, highlighting the effectiveness of the quadratic regression approach in predicting the distance variable. For the 7-parameter model, stability is reached around 1,000 training samples, where both accuracies consistently approach 99%. In contrast, the 8-parameter model, which includes the additional parameter $\Omega_{\Lambda_{2}}$, requires a significantly larger dataset to stabilize. Stability is achieved at approximately 2,500 training samples, with both accuracies converging near 99%. This extended stabilization period reflects the added complexity introduced by the extra parameter, which necessitates a larger dataset for effective generalization. Despite the inclusion of $\Omega_{\Lambda_{2}}$, key parameters such as $A_{s}$ and $\omega_{b}$ remain highly stable across both models, demonstrating their robustness to changes in model complexity.
Minimum distance prediction and check
The regression results show that the 7-parameter and 8-parameter models predict minimum distance values of 27.88 and 24.39 (both rounded), respectively; see Table 2. We again note that the sign of $\Omega_{\Lambda_{2}}$ is negative; the same sign is observed in [27], where its physical significance is discussed. While the 8-parameter model achieves a lower predicted minimum distance, the inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model led to adjustments in other feature values, most notably $\Omega_{\Lambda_{3}}$, which influenced the predicted distance outcome. To validate these predictions, we calculated the distances for the corresponding parameters values; the actual CLASS-computed distance values turn out to be
distance of 7-parameter model distance of 8-parameter model
26.03 24.95
where the values have been rounded. This result confirms that the 8-parameter model achieves a slightly lower minimum distance.
Variable 7-Variable Model 8-Variable Model
$\Omega_{\Lambda}$ - 2.077437E-08
$\Omega_{\Lambda}/\Omega_{\Lambda 3}$ 0.00233471 0.00432285
$h$ 0.67733813 0.67537001
$\omega_{b}$ 0.02238564 0.02238631
$\omega_{cdm}$ 0.11984093 0.11985634
$A_{s}$ 2.094146E-09 2.09484E-09
$n_{s}$ 0.96599393 0.96583363
$\tau_{reio}$ 0.05295666 0.0530449
predicted distance 27.877821 24.391195
Table 2: Optimized feature values for minimum distance in 7- and 8- parameter models
3.2.2 The results with bootstrap resampling
Variable 7-Variable Model (Mean [95% CI]) 8-Variable Model (Mean [95% CI])
$\Omega_{\Lambda_{2}}$ -2.070386E-08 [-2.073444E-08, -2.068444E-08]
$\Omega_{K}/\Omega_{\Lambda_{3}}$ 2.336606E-03 [2.331111E-03, 2.338889E-03] 4.282391E-03 [4.2700E-03, 4.304444E-03]
$h$ 0.6773487 [0.6773333, 0.6780] 0.6753918 [0.6753333, 0.6755]
$\omega_{b}$ 0.02239074 [0.02238667, 0.02239778] 0.02239451 [0.02239444, 0.02239444]
$\omega_{cdm}$ 0.1198824 [0.1196667, 0.1198889] 0.1198889 [0.1198889, 0.1198889]
$\Lambda_{s}$ 2.098202E-09 [2.096667E-09, 2.101111E-09] 2.099984E-09 [2.098889E-09, 2.103333E-09]
$n_{s}$ $0.9660145$ [$0.9660$, $0.9665$] $0.9659412$ [$0.9658889$, $0.9660$]
$\tau_{reid}$ $0.05385517$ [$0.05350$, $0.0548333$] $0.05422211$ [$0.05400$, $0.05500$]
predicted distance $33.30075$ [$33.08834$, $33.6667$] $33.35167$ [$33.06932$, $33.69121$]
Table 3: Optimized Feature Values and 95% Confidence Intervals
To rigorously assess the variability and robustness of the regression-based optimization process, we incorporate a bootstrap resampling procedure - a widely used non-parametric technique in statistical analysis. This approach allows us to empirically quantify the uncertainty in our best-fit cosmological parameters and predicted distance values, independent of any prior distributional assumptions.
In this analysis, we perform 1,000 bootstrap resampling iterations for both the 7-parameter and 8-parameter datasets. In each iteration, the training set (comprising 70% of the full dataset, i.e., 3,500 samples) is resampled with replacement to generate a new training subset of the same size. The quartic polynomial regression model—identified earlier as the best-performing model based on $R^{2}$ and mean squared error (MSE)—is then refitted using the resampled data. For each refitted model, we identify the parameter configuration that minimizes the predicted distance, effectively simulating the optimization process under perturbed training data conditions. This repeated process yields a distribution of best-fit parameter values and predicted distance outcomes across the 1,000 bootstrap samples.
From these distributions, we compute empirical 95% confidence intervals (CIs) by extracting the 2.5th and 97.5th percentiles of each parameter’s distribution. These intervals provide a direct, data-driven estimate of the sampling variability inherent in our modeling framework. Unlike Bayesian credible intervals, which require specification of prior distributions, bootstrap-derived CIs rely solely on the observed data and model fit, making them especially suitable for complex, high-dimensional systems where formal likelihood functions may be intractable or computationally expensive.
The results of this analysis, summarized in Table 3, reveal consistently narrow confidence intervals for most cosmological parameters in both the 7- and 8-parameter models. This suggests a high degree of robustness in the regression-based optimization process: the estimated values are relatively stable across random perturbations of the input data. The predicted distance values likewise exhibit limited variability under bootstrap resampling.
Interestingly, the inclusion of the additional parameter $\Omega_{\Lambda_{2}}$ in the 8-parameter model introduces modest shifts in the distributions of several other parameters - most notably the Hubble parameter $h$, the scalar spectral index $n_{s}$, and the reionization optical depth $\tau_{reio}$. These shifts are accompanied by slightly broader confidence intervals and increased standard deviation in the predicted distance values, indicating a mild increase in model complexity and sensitivity to input variability. Nevertheless, the confidence intervals for these parameters in both models overlap substantially, suggesting that the two regression models produce statistically consistent outputs.
Taken together, the bootstrap results lend strong support to the generalizability and robustness of our quartic regression models. They confirm that the optimized parameter values obtained via our data-driven procedure are not artifacts of a particular sample realization, but instead reflect stable features of the underlying parameter-distance relationship as encoded in the CLASSy-generated datasets.
3.3 (Un)naturalness of the order of $\Omega_{\Lambda_{2}}$
Let us now consider the naturalness of the order of magnitude of $\Omega_{\Lambda_{2}}$, which has been found to be on the order of $10^{-8}$. The solution to the CC problem proposed in [5, 6, 7], which was briefly reviewed in section 2, suggests that $\Omega_{\Lambda_{2}}$ should be related to $\Omega_{\Lambda_{1}}$ through a factor of $10^{-(2\sim 3)}$, i.e., $\Omega_{\Lambda_{2}}\sim\Omega_{\Lambda_{1}}\times 10^{-(2\sim 3)}$. Given that $\Omega_{\Lambda_{1}}\simeq 0.7$, the natural expectation for the order of $\Omega_{\Lambda_{2}}$ is $\Omega_{\Lambda_{2}}\simeq 10^{-(3\sim 4)}$. However, the numerical analysis carried out in the present study yields a value of $\Omega_{\Lambda_{2}}\sim 10^{-8}$, which is several orders of magnitude smaller than the expected range of $10^{-(3\sim 4)}$.
This discrepancy, where the order of $10^{-8}$ seems much smaller than the expected $10^{-(3\sim 4)}$, raises the question of whether this result is unnatural, and if so, whether it can be explained. It would indeed be beneficial to explore this gap further and attempt to provide at least some level of explanation for it. However, before delving into a more detailed analysis, let us briefly recall the nature of the proposal in [5, 6, 7] and the challenges it presents.
The solution of the CC problem suggested in [5, 6, 7] is not without its ”cost.” This cost, however, is not unique to the proposal at hand, but rather reflects a more general phenomenon inherent in renormalization schemes. The approach used in those works involves an appropriate renormalization scheme, and as such, it is not free from a certain degree of unnaturalness when applied to other problems. As well known, physical quantities must be independent of the renormalization schemes used. There are infinitely many renormalization schemes that differ by finite renormalization. It is possible, and often necessary, to change the renormalization scheme when an initial choice made for one problem becomes inconvenient for another.
In the case of the CC problem, this unnaturalness, which is nothing but the well-known fine-tuning, arises specifically due to the onshell renormalization scheme. This scheme uses the physical value of certain parameters, such as the Higgs mass, directly, which leads to a contribution to the CC that is much larger than the observed value. This prompts an important question: if using the physical mass in the onshell scheme leads to this issue, why not consider an alternative renormalization scheme that could potentially avoid or mitigate the problem? See [19] (and references therein) for a recent work with similar ideas.)
In this spirit, the proposal in [5, 6] suggests an alternative approach: rather than using the physical masses directly, one could adopt renormalized mass values that are sufficiently small to avoid causing fine-tuning issues. More specifically, it is suggested that mass values on the order of the temperature might be more appropriate. This would ideally lead to a solution that is less prone to the unnaturalness arising from the onshell renormalization scheme.
The current situation, despite appearing somewhat unnatural with $\Omega_{\Lambda_{2}}\sim 10^{-8}$, may not necessarily represent an insurmountable obstacle. Instead, it could reflect a situation where the issue might be resolved by adopting a different renormalization scheme. The key feature of this alternative scheme is that the masses are adjusted in such a way that the mass-dependent terms in the potential of eq. (22), which were neglected in the previous renormalization scheme applied to the CC problem, are kept small but not negligibly so. This should lead to a slower convergence in the large-temperature expansion since the order of $\frac{m}{T}$, where $m$ denotes the value of the renormalized mass of, say, a Higgs particle, is larger than before. This slow convergence, which was also expected in the proposed CC resolution in [5, 6], could, in turn, suggest that the renormalized cosmological constant $\Lambda_{ren}$ should be set to a value smaller than the observed cosmological constant $\Lambda_{obs}$, by several orders of magnitude. This contrasts with the approach used in the CC problem resolution, where the renormalized and observed values were taken to be on the same order.
Such an adjustment would allow for a potential compensation of the deficit through higher-order terms that slowly converge. As a result, the observed value of $\Omega_{\Lambda_{2}}\sim 10^{-8}$ might not appear as unnatural after all. It is conceivable that higher-order terms in the expansion could shift the one-loop cosmological constant to a value closer to the observed value, thus bridging the gap between the theoretical estimate and the actual observed data.
The situation could also unfold differently on the numerical side. There is a possibility that by repeating the numerical analysis in section 3, and including higher-order terms in the calculation, the optimal value of $\Omega_{\Lambda_{2}}$ could increase from $\sim 10^{-8}$. In this case, the inclusion of these terms might lead to a larger value for $\Omega_{\Lambda_{2}}$, which could help reconcile the numerical result with expectations based on the proposed renormalization scheme. This speculation suggests that the value of $\Omega_{\Lambda_{2}}$ might not be fixed and could evolve as higher-order corrections are included, offering a potential resolution to the discrepancy.
In conclusion, while the order $10^{-8}$ might initially seem unnatural, it is possible that the issue could be addressed through the use of an alternative renormalization scheme that adjusts the mass parameters in a way that makes the solution more natural, or that higher-order terms could lead to a more reasonable outcome. Further analysis is required to clarify the situation, and we will address these possibilities in the conclusion.
4 Conclusion
[…]
While the results presented here reinforce the utility of finite-$T$ QFT effects in cosmological studies, they also open up avenues for future research. A natural next step would be to repeat the full Bayesian parameter estimation using likelihood functions from the Planck 2018 analysis. This more standard approach to parameter inference - though computationally intensive - would provide an important cross-check of our results and help quantify uncertainties more systematically. As noted in the main body, this analysis was not pursued in the present work due to its high computational cost, which would have required access to more powerful computing resources.
One potential improvement can be seen from the residual plots in Figs. 4 and 5, which display certain patterns for both models, particularly for the 8-variable model. While the residuals remain centered around zero, the spread indicates minor deviations from ideal predictions, especially for higher distance values. These residual patterns suggest that while the models are highly accurate, there is room for improvement in capturing some higher-order non-linearities or interactions present in the data. Various different machine learning techniques, including artificial neural networks, have been and are currently being tried. So far, the quartic regression and artificial neural network methods turn out to be best-performing. The result of the artificial neural network [27] shares the same qualitative features as those obtained here. Some complementary aspects, such as feature importance, are reported therein.
The higher-order terms may well play important roles - an expectation motivated by the slow convergence observed in previous studies of the cosmological constant problem [5, 6]. Refining the current approach to incorporate these higher-order corrections, as well as extending machine learning methods for enhanced predictive accuracy, will thus be important next steps. In particular, it will be worth including an additional parameter associated with the linear temperature term, $\sim\frac{1}{a}$, and examining its impact on the rest of the parameters. [^12] The ellipsis in eq. (24) includes both higher-order loop corrections and high-order terms in the $\frac{m}{T}$ expansion. Analyzing the former terms would require significantly greater effort. While our present numerical study indicates that leading-order finite-temperature effects alone do not resolve the tension, it remains to be seen whether this further extended model could potentially ameliorate it. In contrast, see, e.g., [28] for a study in which a novel renormalization scheme achieves a similar goal. Given the exponential growth in the complexity of the parameter space, addressing these higher-order effects and thoroughly exploring their implications - especially with more complete Bayesian methods - will likely require supercomputing resources. We will report progress on these issues elsewhere.
Appendix A Onshell thermodynamics vs. offshell 1PI analysis
In section 2.2, a potential redundancy of $\Omega_{\Lambda_{2}}$ (and $\Omega_{\Lambda_{3}}$) was addressed. Nonetheless, a lingering question may ramain, given that thermodynamic relations, such as $p=\frac{1}{3}\rho$, can also be derived in a quantum field-theoretic thermodynamic framework by introducing the Helmholtz free energy. Such a framework, while useful in certain contexts, is not entirely suitable for the present discussion, as we will explain.
First, let us briefly review the framework. Consider again the Einstein-scalar system in (21):
$S_{\text{ren}}=-\int d^{4}x\left[\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\right].$ (A.1)
We define
$V_{\text{ren}}\equiv\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4},\quad m_{\phi}^{2}\equiv m^{2}+\frac{1}{2}\lambda\phi^{2}.$ (A.2)
The one-loop effective potential at finite temperature can be written as (see, e.g., [29] and the review in [4])
$V_{\text{one-loop}}(\phi,T)=V_{\text{ren}}+\hbar J(m_{\phi},T),\quad\text{where}\quad J(m_{\phi},T)=\int\sum_{n}\frac{1}{2}\ln(K^{2}+m_{\phi}^{2}).$ (A.3)
This leads to the expression
$J(m_{\phi},T)=-\frac{\pi^{2}}{90}T^{4}\!+\!\frac{m_{\phi}^{2}}{24}T^{2}\!-\!\frac{1}{12\pi}m_{\phi}^{3}T\!-\!\frac{m_{\phi}^{4}}{2(4\pi)^{2}}\left[\ln\left(\frac{\bar{\mu}e^{\gamma}}{4\pi T}\right)\!+\!\frac{1}{2\epsilon}\right]\!+\!\frac{\zeta(3)}{3(4\pi)^{4}}\frac{m_{\phi}^{6}}{T^{2}}+\dots$ (A.4)
where $\ln\mu^{2}=\ln(4\pi\mu^{2})-\gamma$ is the renormalization scale, $\gamma$ is the Euler constant, $\epsilon$ is the regularization parameter, and $\zeta(3)$ is the Riemann zeta function. Note that $\hbar$ is explicitly displayed. In the framework under consideration, $V_{\text{one-loop}}(\phi,T)$ is identified as the Helmholtz free energy density, and thermodynamic relations such as $p=\frac{1}{3}\rho$ can be derived for a massless case by setting the free energy density $f=\frac{\pi^{2}}{90}T^{4}$.
However, there are several important issues with applying this derivation directly to the cosmological context:
  1. 1.
    Quantum vs. Classical Thermodynamics: As seen in (A.3), $J(m_{\phi},T)$ represents a quantum correction, in contrast to the classical kinetic theory derivation of thermodynamic relations such as $p=\frac{1}{3}\rho$. The kinetic theory treatment is purely classical and does not involve quantum corrections.
  2. 2.
    On-shell Condition: For the derivation of $p=\frac{1}{3}\rho$, the scalar field $\phi$ must be set to zero, i.e., $\phi=0$, which is a trivial solution of the field equation. This implies an on-shell analysis, where the field $\phi$ is fixed to a solution. However, this is not the objective of the present work. The current analysis keeps the $\phi$-dependent part in order to obtain the quantum-corrected offshell aciton, namely the 1PI action, and focuses on the $\phi$-independent contributions to the cosmological constant. Thus, the field-theoretic derivation of $p=\frac{1}{3}\rho$ in the quantum field-theoretic thermodynamics does not imply that, for example, the $T^{4}$ contribution from vacuum diagrams is redundant to the radiation contribution.
  3. 3.
    Massive Particles in the Standard Model: In the Standard Model (SM), there are massive particles that cannot be regarded as radiation. These particles contribute to the $T^{4}$- and $T^{2}$-terms in the cosmological constant, even though they do not fit within the classical radiation picture (i.e., $p=\frac{1}{3}\rho$).
[…]
On the other hand, for matter components like dust (e.g., stars, galaxies), a purely classical treatment is appropriate. These objects are not described by Feynman diagrams but rather by kinetic theory. Hence, the only sensible approach to describing them is via kinetic theory or hydrodynamics. Since it is desirable to treat all matter components uniformly, a classical kinetic theory treatment of all matter components is the appropriate setup.
To further clarify the distinction, consider a slightly different, yet conceptually simpler approach. This approach, while arriving at the same starting point as before within the typical approximations used in cosmology, begins by formulating an effective field theory for gravity and SM particles, without including the kinetic theory sector. (The term ”effective” refers here to an effective field theory derived by keeping only the leading terms in the derivative expansion of the 1PI action.) This effective action accounts for quantum corrections, including a $T^{4}$ term (and a $T^{2}$ term) in the CC sector. Subsequently, a kinetic or hydrodynamic system - which includes matter compoenents such as $\rho_{r}$ or $\rho_{cdm}$ - can be coupled to this effective field theory. For a purely hydrodynamic system, all SM fields can be set to zero.
Appendix B Statistical glossary
Regression:
A statistical technique used to model and analyze the relationship between a dependent variable and one or more independent variables (also called features or predictors). The goal is to estimate how changes in the independent variables influence the dependent variable. Common types include linear regression, where the relationship is modeled as a straight line, and polynomial regression, where higher-degree terms allow modeling of non-linear patterns.
$R^{2}$ (Coefficient of determination):
A statistical measure that represents the proportion of variance in the dependent variable that is predictable from the independent variables. It ranges from 0 to 1, with values closer to 1 indicating better model performance.
MSE (Mean squared error):
A measure of the average squared difference between the observed actual outcomes and the outcomes predicted by the model. It is given by
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$
where $y_{i}$ is the actual value, $\hat{y}_{i}$ is the predicted value, and $n$ is the number of observations.
Standardization:
A preprocessing technique used to scale features so they have a mean of 0 and a standard deviation of 1. This ensures that features contribute equally to the model, especially important for distance-based and regularized methods.
Residuals:
The difference between observed values and the values predicted by a regression model. For a given observation, the residual is $e_{i}=y_{i}-\hat{y}_{i}$. Residuals are used to diagnose the fit of a regression model.
Learning curve:
A plot showing how a model’s performance (typically accuracy or error) on the training and validation datasets changes as the training set size increases. It helps assess whether adding more data improves model performance.
Overfitting:
A modeling error that occurs when a regression model is too complex and captures noise rather than the underlying pattern in the data. It usually results in high training accuracy but poor generalization to new data.
Cross-validation:
A statistical method used to estimate the performance of a model on an independent dataset. In $k$-fold cross-validation, the dataset is divided into $k$ subsets; the model is trained on k1k-1 of them and tested on the remaining subset. This process is repeated $k$ times, each time with a different test set.